Same case as above: Will Rogers owns and manages a taco and soft-drink stand in Lubbock near Texas Tech University. On Saturdays, his busiest day, he can service 30 customers per hour on the average but he only gets 20 customers per hour on average. Because Will can wait on 50% more customers than actually visits his restaurant, it does not make sense to him that he should have any waiting lines. After looking at the problem you find that it follows the six conditions for a single-server waiting line. What is the average number of customers in the queue?
Answer #1

Arrival rate = λ = 20 /hour

Service Rate = μ = 30/hour

The average number of customers in the queue = λ2/μ(μ-λ) = 202/(30*(30-20)) = 1.33